OpenStudy (anonymous):

In the system shown below, what is the sum of all of the x-coordinates of all solutions? x^2+4y^2=100 4y-x^2=-20 please please explain I am very stuck

geerky42 (geerky42):

Try to use elimination method: \[ ~~~~x^2+4y^2=100 \\ \underline{+~-x^2+4y=-20}\\~~~~~~~~~~~~~~?\]

geerky42 (geerky42):

add equations together, you should eliminate \(x^2\), leaves you \(4y^2+4y=80\)

OpenStudy (aum):

They want the sum of all the x-coordinates of the solution. Eliminate y. Find y from the second equation, substitute it in y and you will have an equation in x.

geerky42 (geerky42):

Explain how to eliminate y please @aum

OpenStudy (aum):

4y-x^2=-20 add x^2 4y = x^2-20 divide by 4 y = (x^2-20)/4 Put this in the first equation and simplify.

geerky42 (geerky42):

it's more of a substitution than elimination, but that's good move though

OpenStudy (anonymous):

I put (x^2-20)/4 back into the equation and my answer was bizarre

geerky42 (geerky42):

I know, right lol... So you have \(x^2+4\left(\dfrac{x^2-20}{4}\right)^2=100\) You can simplify \(\left(\dfrac{x^2-20}{4}\right)^2\) to \(\dfrac{(x^2-20)^2}{4^2}\) What is \((x^2-20)^2\) ?

OpenStudy (anonymous):

x^4-40x^2-400

geerky42 (geerky42):

Almost. it's \(\Large x^4-40x^2+400\)

OpenStudy (anonymous):

sorry typo haha

OpenStudy (anonymous):

so I multiplied it out and got 4x^4-160x^2+1600

OpenStudy (anonymous):

so it would be 4x^4-158x^2+1600=100?

geerky42 (geerky42):

lol okay, so you now have \(x^2+4\cdot\dfrac{x^4-40x^2+400}{4^2}=100 \\ x^2+\cancel{4}\cdot\dfrac{x^4-40x^2+400}{4^{\cancel2}}=100\\x^2+\dfrac{x^4-40x^2+400}{4}=100\) Hey, you hate denominator, right? me too, so let multiply both sides by 4 to get \(4x^2+x^4-40x^2+400=400\)

OpenStudy (anonymous):

wow im way off, so then we solve for x okay

geerky42 (geerky42):

do you understand what I just did, right?

OpenStudy (anonymous):

Yes, I see it now

geerky42 (geerky42):

ok great, so we now have \(4x^2+x^4-40x^2+400=400 \\\rightarrow x^4-36x^2+400=400\\\rightarrow x^4-36x^2=0\) Do you know what to do next?

OpenStudy (anonymous):

take out a greatest common factor (x^2)

OpenStudy (anonymous):

so the answers are x=6 x=-6?

geerky42 (geerky42):

and x=0 too. Question wants sum of all of the x-coordinates of all solutions So you have -6+0+6=?

OpenStudy (anonymous):

So zero

geerky42 (geerky42):

right. your answer is 0

OpenStudy (anonymous):

thank you so much!

geerky42 (geerky42):

love how lengthy and harsh problem ended up get us 0 haha

geerky42 (geerky42):

welcome!

OpenStudy (anonymous):

I know right